# A Continuous-Time Ensemble Kalman–Bucy Smoother for Causal Inference and Model Discovery

### 28 Apr 2026 v1

@@ 2.2 The Forward Pass: EnKBF
We present a stochastic variant of EnKBF in this subsection, dropping the subscripts $\mathrm{f}$ for notational simplicity. First, assume that the observation drift $\mathbf{h}$ is differentiable with respect to $\mathbf{x}$. Let $\mathbf{h}\approx\mathbf{H}\mathbf{x}$ denote its local linearization, defined through its Jacobian $\mathbf{H}$. The classical discrete-time Kalman filter under the lens of Bayesian inference consists of two consecutive steps [25, 32, 34]: the update or forecast step, which propagates the dynamics to obtain a model-based prior, and the analysis step, which assimilates new observations into the prior, through the likelihood, to produce a posterior via Bayes’ theorem. Translating this viewpoint to the continuous-time setting, we regard the analysis step as incorporating observations via a continuous flow rather than abrupt discrete updates. This is formulated by introducing a fictitious assimilation time interval $s\in[0,1]$ at each physical time $t$ [47, 57, 58]. Specifically, for each ensemble member, we generate a simulated observation increment
$\mathrm{d}\mathbf{y}^{(i)}:=\mathbf{H}\mathbf{x}^{(i)}\,\mathrm{d}t+\mathbf{\Gamma}^{1/2}\,\mathrm{d}\mathbf{W}^{(i)},$ (2)
@@
where $\{\mathbf{W}^{(i)}\}_{i=1,\ldots,m}$ are mutually independent standard Wiener processes. At each physical time $t$, the analysis step update for the $i$-th ensemble member, $i=1,\dots,m$, is then governed by the fictitious-time SDE
$\mathrm{d}\mathbf{x}^{(i)}=\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\left(\,\mathrm{d}\mathbf{y}-\,\mathrm{d}\mathbf{y}^{(i)}\right),\qquad s\in[0,1],$ (3)
@@
with Kalman gain $\mathbf{K}:=\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}$. Adding (3) as an additional drift to the “forecast” dynamics (1a) yields the stochastic continuous-time EnKBF formulation with a linear observation operator $\mathbf{H}$ [57, 59]:
$\mathrm{d}\mathbf{x}^{(i)}=\mathbf{f}(\mathbf{x}^{(i)},t)\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}+\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\left(\,\mathrm{d}\mathbf{y}-\,\mathrm{d}\mathbf{y}^{(i)}\right).$ (4)
@@
Treating the simulated observations as random variables corresponding to each ensemble member ensures statistical consistency with the Riccati covariance equation in classical Kalman filter theory; in the absence of the noise term $\mathbf{\Gamma}^{1/2}\,\mathrm{d}\mathbf{W}^{(i)}$ in (2), the filtering covariance would be systematically underestimated [39, 60].
@@
In practical applications, the observation drift $\mathbf{h}$ is often nonlinear. While we rely on the local linearization $\mathbf{h}\approx\mathbf{H}\mathbf{x}$ to derive (4), we apply the following approximations to obtain a derivative-free framework:
$\mathbf{H}\mathbf{x}^{(i)}\approx\mathbf{h}(\mathbf{x}^{(i)},t),$ (5)
$\mathbf{P}\mathbf{H}^{\top}\approx\mathbf{P}_{\mathbf{x}\mathbf{h}}.$ (6)
@@
Here, $\mathbf{P}_{\mathbf{x}\mathbf{h}}$ denotes the cross-covariance matrix between the ensemble and its observation-operator-evolved counterpart:
$\mathbf{P}_{\mathbf{x}\mathbf{h}}:=\frac{1}{m-1}\sum_{i=1}^{m}\left(\mathbf{x}^{(i)}-\overline{\mathbf{x}}\right)\left(\mathbf{h}(\mathbf{x}^{(i)})-\overline{\mathbf{h}}\right)^{\top},$
@@
$\overline{\mathbf{h}}(t):=\frac{1}{m}\sum_{i=1}^{m}\mathbf{h}(\mathbf{x}^{(i)},t).$
Substituting (5)–(6) into (4) yields the practical derivative-free EnKBF formulation for $i=1,\dots,m$:
$\mathrm{d}\mathbf{x}^{(i)}=\mathbf{f}(\mathbf{x}^{(i)},t)\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}+\mathbf{P}_{\mathbf{x}\mathbf{h}}\mathbf{\Gamma}^{-1}\left(\,\mathrm{d}\mathbf{y}-\,\mathrm{d}\mathbf{y}^{(i)}\right),$ (7)
where, under the approximation (5), the simulated observation in (2) becomes $\mathrm{d}\mathbf{y}^{(i)}=\mathbf{h}(\mathbf{x}^{(i)},t)\,\mathrm{d}t+\mathbf{\Gamma}^{1/2}\,\mathrm{d}\mathbf{W}^{(i)}$. We remark that applying (5)–(6) avoids explicit tangent-linear or adjoint computations related to $\mathbf{H}$ [12, 40]. The key assumption behind (6) is
$\mathbf{h}({\mathbf{x}}^{(i)})-\overline{\mathbf{h}}\approx\mathbf{H}\big({\mathbf{x}}^{(i)}-\overline{\mathbf{x}}\big),$
which can be interpreted as a finite-difference approximation of the drift $\mathbf{h}$ when the $i$-th ensemble spread, ${\mathbf{x}}^{(i)}-\overline{\mathbf{x}}$, is sufficiently small.
@@ 2.3 The Backward Pass: EnKBS
Given a filtering ensemble $\mathbf{x}_{\mathrm{f}}^{(i)}(t),i=1,\dots,m,$ produced by the forward stochastic EnKBF (7), the smoother integrates backward from $t=T$ to $t=0$ to refine the ensemble statistics over $[0,T]$ in the context of the full observed trajectory $\mathbf{y}_{[0,T]}$. Denote the smoothing ensemble by $\mathbf{x}_{\mathrm{s}}^{(i)}(t),i=1,\dots,m$. Initializing the backward pass at $t=T$ by
$\mathbf{x}_{\mathrm{s}}^{(i)}(T)=\mathbf{x}_{\mathrm{f}}^{(i)}(T),$ (8)
@@
$\overleftarrow{\,\mathrm{d}}\!\mathbf{x}_{\mathrm{s}}^{(i)}=\mathbf{f}\big(\mathbf{x}_{\mathrm{s}}^{(i)},t\big)\overleftarrow{\,\mathrm{d}}\!t+\mathbf{\Sigma}^{1/2}\overleftarrow{\,\mathrm{d}}\!\mathbf{B}^{(i)}+\mathbf{\Sigma}\mathbf{P}_{\mathrm{f}}^{-1}\big(\mathbf{x}_{\mathrm{s}}^{(i)}-\mathbf{x}_{\mathrm{f}}^{(i)}\big)\overleftarrow{\,\mathrm{d}}\!t,$ (9)
@@
where $\mathbf{P}_{\mathrm{f}}$ is the empirical covariance of filtering ensemble $\mathbf{x}_{\mathrm{f}}^{(i)}$, and $\overleftarrow{\,\mathrm{d}}(\cdot)$ denotes the backward differential that defines a backward Itô integral, i.e., $\overleftarrow{\,\mathrm{d}}(\cdot)=-\,\mathrm{d}(\cdot)$ [52, 61]. In (9), we use the same Wiener process paths $\mathbf{B}^{(i)}$ as in the forward filter (7), as justified by the derivations in A.
@@
To clarify the backward-in-time differential $\overleftarrow{\,\mathrm{d}}(\cdot)$ in (9), we state the Euler–Maruyama time discretization that we use in the practical algorithm. Let $\tau>0$ denote the time step and $t_{k}=k\tau$. We use the subscript $(\cdot)_{k}$ to denote evaluation at time $t_{k}$, e.g., $\mathbf{x}_{k}=\mathbf{x}(t_{k})$. The forward EnKBF (7) is discretized as
@@
$\mathbf{x}_{\mathrm{f},k+1}^{(i)}=\mathbf{x}_{\mathrm{f},k}^{(i)}+\tau\mathbf{f}\left(\mathbf{x}_{\mathrm{f},k}^{(i)},t_{k}\right)+\sqrt{\tau}\mathbf{\Sigma}^{1/2}\mathbf{B}_{k}^{(i)}+\mathbf{P}_{\mathbf{x}\mathbf{h},k}\mathbf{\Gamma}^{-1}\left(\Delta\mathbf{y}_{k+1}-\Delta\mathbf{y}_{k+1}^{(i)}\right),$ (10)
@@
for $k=0,1,\dots,K-1$, where $\mathbf{B}_{k}^{(i)},\mathbf{W}_{k}^{(i)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ are independent, $\Delta\mathbf{y}_{k+1}=\mathbf{y}_{k+1}-\mathbf{y}_{k}$ is the observation increment, and
$\Delta\mathbf{y}_{k+1}^{(i)}:=\tau\mathbf{h}\left(\mathbf{x}_{\mathrm{f},k}^{(i)},t_{k}\right)+\sqrt{\tau}\mathbf{\Gamma}^{1/2}\mathbf{W}_{k}^{(i)}$
@@
is the simulated observation increment for the $i$-th member, consistent with the continuous-time definition (2). The backward EnKBS (9) is discretized as
$\mathbf{x}_{\mathrm{s},k}^{(i)}=\mathbf{x}_{\mathrm{s},k+1}^{(i)}-\tau\mathbf{f}\left(\mathbf{x}_{\mathrm{s},k+1}^{(i)},t_{k+1}\right)-\sqrt{\tau}\mathbf{\Sigma}^{1/2}\mathbf{B}_{k}^{(i)}-\tau\mathbf{\Sigma}\mathbf{P}_{\mathrm{f},k+1}^{-1}\left(\mathbf{x}_{\mathrm{s},k+1}^{(i)}-\mathbf{x}_{\mathrm{f},k+1}^{(i)}\right),$ (11)
@@
for $k=K-1,K-2,\dots,0$, with initialization $\mathbf{x}_{\mathrm{s},K}^{(i)}=\mathbf{x}_{\mathrm{f},K}^{(i)}$. Note that the same noise increments $\mathbf{B}_{k}^{(i)}$ from the forward pass are reused in the backward pass.
@@
As a remark, in the EnKBS formulation (9), the backward dynamics do not depend explicitly on the observation increment $\mathrm{d}\mathbf{y}$ as with the EnKBF in (7). The implicit dependence on $\mathbf{y}$ is instead via the filtering ensemble $\mathbf{x}_{\mathrm{f}}^{(i)}(t)$} computed in the forward pass. Under the assumption of independent signal and observation white noises, the information in $\mathbf{y}$ is already encoded in the future state $\mathbf{x}$ and propagates backward to correct the current estimate along the smoother pass. Conditioned on the future $\mathbf{x}$, the Markov property of the system yields that including any observation $\mathbf{y}$ depending on it does not provide additional information in the absence of cross-interacting noise [49, 52, 62].
@@
Detailed derivations of (9) and its formal connection to the discrete-time RTS smoother in the formal limit of a vanishing observation interval are provided in A. In addition, B shows that the proposed EnKBS recovers the exact mean and covariance equation of the continuous-time RTS smoother in the linear-Gaussian setting [33].
@@ 2.4 Localization and Inflation
In high-dimensional data assimilation with state dimension $n$, a small ensemble size $m\ll n$ inevitably incurs sampling error and cannot fully resolve all dynamic modes. Specifically, the full covariance matrix is $n\times n$, whereas the empirical ensemble covariance $\mathbf{P}$ is rank-deficient with rank at most m1m-1, which is often insufficient to fully capture the dominant unstable modes of the system [63]. This rank deficiency may cause instability or divergence of the filter and smoother and induce spurious, nonphysical long-range correlations [40, 48]. In many operational applications, affordable ensemble sizes typically range from $O(10)$ to $O(100)$, which can be much smaller than the dimension of the unstable subspace [12]. Increasing $m$ to that scale is often computationally prohibitive.
Two standard regularization techniques, localization and inflation, are widely used to address these issues and improve stability in high-dimensional systems. These ad hoc strategies, to some extent, help explain the practical success of ensemble methods in applications such as numerical weather prediction and the geosciences [42].
Localization
Covariance localization, also called Schur localization, regularizes the empirical covariance by damping long-range state-space correlations and thereby increases its rank [12, 47]. A standard approach applies a Schur (element-wise) product between the empirical covariance $\mathbf{P}$ and a short-ranged correlation matrix $\mathbf{C}_{\mathrm{loc}}\in\mathbb{R}^{n\times n}$:
@@
By the Schur product theorem [64], $\mathbf{C}_{\mathrm{loc}}\circ\mathbf{P}$ is positive (semi-)definite whenever $\mathbf{C}_{\mathrm{loc}}$ and $\mathbf{P}$ are positive (semi-)definite. A common choice constructs $\mathbf{C}_{\mathrm{loc}}$ from the Gaspari–Cohn fifth-order piecewise rational function $G$ [65],
@@
Accordingly, the localized formulation of the stochastic EnKBF (7) replaces covariance $\mathbf{P}_{\mathbf{x}\mathbf{h}}$ by $\mathbf{C}_{\mathrm{loc,1}}\circ\mathbf{P}_{\mathbf{x}\mathbf{h}}$ [66]:
$\mathrm{d}\mathbf{x}^{(i)}=\mathbf{f}(\mathbf{x}^{(i)},t)\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}+(\mathbf{C}_{\mathrm{loc,1}}\circ\mathbf{P}_{\mathbf{x}\mathbf{h}})\mathbf{\Gamma}^{-1}\left(\,\mathrm{d}\mathbf{y}-\,\mathrm{d}\mathbf{y}^{(i)}\right),$ (13)
and, for consistency, we also apply a localization to the backward EnKBS (9):
$\overleftarrow{\,\mathrm{d}}\!\mathbf{x}_{\mathrm{s}}^{(i)}=\mathbf{f}\big(\mathbf{x}_{\mathrm{s}}^{(i)},t\big)\overleftarrow{\,\mathrm{d}}\!t+\mathbf{\Sigma}^{1/2}\overleftarrow{\,\mathrm{d}}\!\mathbf{B}^{(i)}+\mathbf{\Sigma}(\mathbf{C}_{\mathrm{loc,2}}\circ\mathbf{P}_{\mathrm{f}})^{-1}\big(\mathbf{x}_{\mathrm{s}}^{(i)}-\mathbf{x}_{\mathrm{f}}^{(i)}\big)\overleftarrow{\,\mathrm{d}}\!t.$ (14)
Inflation
With localization, sampling error and spurious correlations are reduced, but still may accumulate over time. A practical solution to this is multiplicative inflation, which rescales ensemble anomalies around the mean by a factor $\delta^{2}>1$ [67]:
$\mathbf{x}^{(i)}\longleftarrow\overline{\mathbf{x}}+\delta\big(\mathbf{x}^{(i)}-\overline{\mathbf{x}}\big),\qquad i=1,\dots,m.$ (15)
The multiplicative inflation (15) aims to compensate for sampling errors that are the indirect consequence of the nonlinearity in the ensemble forecast, which helps to cure the intrinsic error of ensemble filter schemes [12]. In addition, inflation should be applied only for the forward filtering, not the backward smoothing pass. Since sampling errors are already accounted for in the filter, applying inflation in the backward update would result in a suboptimal solution [68].

### 03 May 2026 v2

@@ 1.2 The Forward Pass: EnKBF
We present a stochastic variant of EnKBF in this subsection, dropping the subscripts $\mathrm{f}$ for notational simplicity. First, assume that the observation drift $\vec{h}$ is differentiable with respect to $\vec{x}$. Let $\vec{h}\approx\mathbf{H}\vec{x}$ denote its local linearization, defined through its Jacobian $\mathbf{H}$. The classical discrete-time Kalman filter under the lens of Bayesian inference consists of two consecutive steps (Simon, 2006; Kalman, 1960; Jazwinski, 1970): the update or forecast step, which propagates the dynamics to obtain a model-based prior, and the analysis step, which assimilates new observations into the prior, through the likelihood, to produce a posterior via Bayes’ theorem. Translating this viewpoint to the continuous-time setting, we regard the analysis step as incorporating observations via a continuous flow rather than abrupt discrete updates. This is formulated by introducing a fictitious assimilation time interval $s\in[0,1]$ at each physical time $t$ (Bergemann and Reich, 2010a, 2012, b). Specifically, for each ensemble member, we generate a simulated observation increment
$\mathrm{d}\vec{y}^{(i)}:=\mathbf{H}\vec{x}^{(i)}\mathrm{d}t+\mathbf{\Gamma}^{1/2}\mathrm{d}\vec{W}^{(i)},$ (2)
@@
where $\{\vec{W}^{(i)}\}_{i=1,\ldots,m}$ are mutually independent standard Wiener processes. At each physical time $t$, the analysis step update for the $i$-th ensemble member, $i=1,\dots,m$, is then governed by the fictitious-time SDE
$\mathrm{d}\vec{x}^{(i)}=\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\left(\mathrm{d}\vec{y}-\mathrm{d}\vec{y}^{(i)}\right),\qquad s\in[0,1],$ (3)
@@
with Kalman gain $\mathbf{K}:=\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}$. Adding Eq. (3) as an additional drift to the “forecast” dynamics Eq. (1a) yields the stochastic continuous-time EnKBF formulation with a linear observation operator $\mathbf{H}$ (Bergemann and Reich, 2012; Reich, 2011):
$\mathrm{d}\vec{x}^{(i)}=\vec{f}(\vec{x}^{(i)},t)\mathrm{d}t+\mathbf{\Sigma}^{1/2}\mathrm{d}\vec{B}^{(i)}+\mathbf{P}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\left(\mathrm{d}\vec{y}-\mathrm{d}\vec{y}^{(i)}\right).$ (4)
@@
Treating the simulated observations as random variables corresponding to each ensemble member ensures statistical consistency with the Riccati covariance equation in classical Kalman filter theory; in the absence of the noise term $\mathbf{\Gamma}^{1/2}\mathrm{d}\vec{W}^{(i)}$ in Eq. (2), the filtering covariance would be systematically underestimated (Burgers et al., 1998; Sakov and Oke, 2008).
@@
In practical applications, the observation drift $\vec{h}$ is often nonlinear. While we rely on the local linearization $\vec{h}\approx\mathbf{H}\vec{x}$ to derive Eq. (4), we apply the following approximations to obtain a derivative-free framework:
$\mathbf{H}\vec{x}^{(i)}\approx\vec{h}(\vec{x}^{(i)},t),$ (5)
$\mathbf{P}\mathbf{H}^{\top}\approx\mathbf{P}_{\vec{x}\vec{h}}.$ (6)
@@
Here, $\mathbf{P}_{\vec{x}\vec{h}}$ denotes the cross-covariance matrix between the ensemble and its observation-operator-evolved counterpart:
$\mathbf{P}_{\vec{x}\vec{h}}:=\frac{1}{m-1}\sum_{i=1}^{m}\left(\vec{x}^{(i)}-\overline{\vec{x}}\right)\left(\vec{h}(\vec{x}^{(i)})-\overline{\vec{h}}\right)^{\top},$
@@
$\overline{\vec{h}}(t):=\frac{1}{m}\sum_{i=1}^{m}\vec{h}(\vec{x}^{(i)},t).$
Substituting Eq. (5)–Eq. (6) into Eq. (4) yields the practical derivative-free EnKBF formulation for $i=1,\dots,m$:
$\mathrm{d}\vec{x}^{(i)}=\vec{f}(\vec{x}^{(i)},t)\mathrm{d}t+\mathbf{\Sigma}^{1/2}\mathrm{d}\vec{B}^{(i)}+\mathbf{P}_{\vec{x}\vec{h}}\mathbf{\Gamma}^{-1}\left(\mathrm{d}\vec{y}-\mathrm{d}\vec{y}^{(i)}\right),$ (7)
where, under the approximation (5), the simulated observation in Eq. (2) becomes $\mathrm{d}\vec{y}^{(i)}=\vec{h}(\vec{x}^{(i)},t)\mathrm{d}t+\mathbf{\Gamma}^{1/2}\mathrm{d}\vec{W}^{(i)}$. We remark that applying Eq. (5)–Eq. (6) avoids explicit tangent-linear or adjoint computations related to $\mathbf{H}$ (Asch et al., 2016; Evensen, 2003). The key assumption behind Eq. (6) is
$\vec{h}({\vec{x}}^{(i)})-\overline{\vec{h}}\approx\mathbf{H}\big({\vec{x}}^{(i)}-\overline{\vec{x}}\big),$
which can be interpreted as a finite-difference approximation of the drift $\vec{h}$ when the $i$-th ensemble spread, ${\vec{x}}^{(i)}-\overline{\vec{x}}$, is sufficiently small.
@@ 1.3 The Backward Pass: EnKBS
Given a filtering ensemble $\vec{x}_{\mathrm{f}}^{(i)}(t),i=1,\dots,m,$ produced by the forward stochastic EnKBF Eq. (7), the smoother integrates backward from $t=T$ to $t=0$ to refine the ensemble statistics over $[0,T]$ in the context of the full observed trajectory $\vec{y}_{[0,T]}$. Denote the smoothing ensemble by $\vec{x}_{\mathrm{s}}^{(i)}(t),i=1,\dots,m$. Initializing the backward pass at $t=T$ by
$\vec{x}_{\mathrm{s}}^{(i)}(T)=\vec{x}_{\mathrm{f}}^{(i)}(T),$ (8)
@@
$\overleftarrow{\mathrm{d}}\!\vec{x}_{\mathrm{s}}^{(i)}=\vec{f}\big(\vec{x}_{\mathrm{s}}^{(i)},t\big)\overleftarrow{\mathrm{d}}\!t+\mathbf{\Sigma}^{1/2}\overleftarrow{\mathrm{d}}\!\vec{B}^{(i)}+\mathbf{\Sigma}\mathbf{P}_{\mathrm{f}}^{-1}\big(\vec{x}_{\mathrm{s}}^{(i)}-\vec{x}_{\mathrm{f}}^{(i)}\big)\overleftarrow{\mathrm{d}}\!t,$ (9)
@@
where $\mathbf{P}_{\mathrm{f}}$ is the empirical covariance of filtering ensemble $\vec{x}_{\mathrm{f}}^{(i)}$, and $\overleftarrow{\mathrm{d}}(\cdot)$ denotes the backward differential that defines a backward Itô integral, i.e., $\overleftarrow{\mathrm{d}}\!(\cdot)=-\mathrm{d}(\cdot)$ (Crisan and Rozovskii, 2011; Rozovsky and Lototsky, 2018). In Eq. (9), we use the same Wiener process paths $\vec{B}^{(i)}$ as in the forward filter Eq. (7), as justified by the derivations in Appendix A.
@@
To clarify the backward-in-time differential $\overleftarrow{\mathrm{d}}\!(\cdot)$ in Eq. (9), we state the Euler–Maruyama time discretization that we use in the practical algorithm. Let $\tau>0$ denote the time step and $t_{k}=k\tau$. We use the subscript $(\cdot)_{k}$ to denote evaluation at time $t_{k}$, e.g., $\vec{x}_{k}=\vec{x}(t_{k})$. The forward EnKBF Eq. (7) is discretized as
@@
$\vec{x}_{\mathrm{f},k+1}^{(i)}=\vec{x}_{\mathrm{f},k}^{(i)}+\tau\vec{f}\left(\vec{x}_{\mathrm{f},k}^{(i)},t_{k}\right)+\sqrt{\tau}\mathbf{\Sigma}^{1/2}\vec{B}_{k}^{(i)}+\mathbf{P}_{\vec{x}\vec{h},k}\mathbf{\Gamma}^{-1}\left(\Delta\vec{y}_{k+1}-\Delta\vec{y}_{k+1}^{(i)}\right),$ (10)
@@
for $k=0,1,\dots,K-1$, where $\vec{B}_{k}^{(i)},\vec{W}_{k}^{(i)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ are independent, $\Delta\vec{y}_{k+1}=\vec{y}_{k+1}-\vec{y}_{k}$ is the observation increment, and
$\Delta\vec{y}_{k+1}^{(i)}:=\tau\vec{h}\left(\vec{x}_{\mathrm{f},k}^{(i)},t_{k}\right)+\sqrt{\tau}\mathbf{\Gamma}^{1/2}\vec{W}_{k}^{(i)}$
@@
is the simulated observation increment for the $i$-th member, consistent with the continuous-time definition Eq. (2). The backward EnKBS Eq. (9) is discretized as
$\vec{x}_{\mathrm{s},k}^{(i)}=\vec{x}_{\mathrm{s},k+1}^{(i)}-\tau\vec{f}\left(\vec{x}_{\mathrm{s},k+1}^{(i)},t_{k+1}\right)-\sqrt{\tau}\mathbf{\Sigma}^{1/2}\vec{B}_{k}^{(i)}-\tau\mathbf{\Sigma}\mathbf{P}_{\mathrm{f},k+1}^{-1}\left(\vec{x}_{\mathrm{s},k+1}^{(i)}-\vec{x}_{\mathrm{f},k+1}^{(i)}\right),$ (11)
@@
for $k=K-1,K-2,\dots,0$, with initialization $\vec{x}_{\mathrm{s},K}^{(i)}=\vec{x}_{\mathrm{f},K}^{(i)}$. Note that the same noise increments $\vec{B}_{k}^{(i)}$ from the forward pass are reused in the backward pass.
@@
As a remark, in the EnKBS formulation Eq. (9), the backward dynamics do not depend explicitly on the observation increment $\mathrm{d}\vec{y}$ as with the EnKBF in Eq. (7). The implicit dependence on $\vec{y}$ is instead via the filtering ensemble $\vec{x}_{\mathrm{f}}^{(i)}(t)$} computed in the forward pass. Under the assumption of independent signal and observation white noises, the information in $\vec{y}$ is already encoded in the future state $\vec{x}$ and propagates backward to correct the current estimate along the smoother pass. Conditioned on the future $\vec{x}$, the Markov property of the system yields that including any observation $\vec{y}$ depending on it does not provide additional information in the absence of cross-interacting noise (Andreou and Chen, 2024a; Crisan and Rozovskii, 2011; Durrett, 2019).
@@
Detailed derivations of Eq. (9) and its formal connection to the discrete-time RTS smoother in the formal limit of a vanishing observation interval are provided in Appendix A. In addition, Appendix B shows that the proposed EnKBS recovers the exact mean and covariance equation of the continuous-time RTS smoother in the linear-Gaussian setting (Rauch et al., 1965).
@@ 1.4 Localization and Inflation
In high-dimensional data assimilation with state dimension $n$, a small ensemble size $m\ll n$ inevitably incurs sampling error and cannot fully resolve all dynamic modes. Specifically, the full covariance matrix is $n\times n$, whereas the empirical ensemble covariance $\mathbf{P}$ is rank-deficient with rank at most m1m-1, which is often insufficient to fully capture the dominant unstable modes of the system (Bocquet and Carrassi, 2017). This rank deficiency may cause instability or divergence of the filter and smoother and induce spurious, nonphysical long-range correlations (Evensen, 2003; Houtekamer and Mitchell, 1998). In many operational applications, affordable ensemble sizes typically range from $O(10)$ to $O(100)$, which can be much smaller than the dimension of the unstable subspace (Asch et al., 2016). Increasing $m$ to that scale is often computationally prohibitive.
Two standard regularization techniques, localization and inflation, are widely used to address these issues and improve stability in high-dimensional systems. These ad hoc strategies, to some extent, help explain the practical success of ensemble methods in applications such as numerical weather prediction and the geosciences (Houtekamer and Zhang, 2016).
Localization. Covariance localization, also called Schur localization, regularizes the empirical covariance by damping long-range state-space correlations and thereby increases its rank (Asch et al., 2016; Bergemann and Reich, 2010a). A standard approach applies a Schur (element-wise) product between the empirical covariance $\mathbf{P}$ and a short-ranged correlation matrix $\mathbf{C}_{\mathrm{loc}}\in\mathbb{R}^{n\times n}$:
@@
By the Schur product theorem (Horn and Johnson, 1985), $\mathbf{C}_{\mathrm{loc}}\circ\mathbf{P}$ is positive (semi-)definite whenever $\mathbf{C}_{\mathrm{loc}}$ and $\mathbf{P}$ are positive (semi-)definite. A common choice constructs $\mathbf{C}_{\mathrm{loc}}$ from the Gaspari–Cohn fifth-order piecewise rational function $G$ (Gaspari and Cohn, 1999),
@@
Accordingly, the localized formulation of the stochastic EnKBF Eq. (7) replaces covariance $\mathbf{P}_{\vec{x}\vec{h}}$ by $\mathbf{C}_{\mathrm{loc,1}}\circ\mathbf{P}_{\vec{x}\vec{h}}$ (Khare et al., 2008):
$\mathrm{d}\vec{x}^{(i)}=\vec{f}(\vec{x}^{(i)},t)\mathrm{d}t+\mathbf{\Sigma}^{1/2}\mathrm{d}\vec{B}^{(i)}+(\mathbf{C}_{\mathrm{loc,1}}\circ\mathbf{P}_{\vec{x}\vec{h}})\mathbf{\Gamma}^{-1}\left(\mathrm{d}\vec{y}-\mathrm{d}\vec{y}^{(i)}\right),$ (13)
and, for consistency, we also apply a localization to the backward EnKBS Eq. (9):
$\overleftarrow{\mathrm{d}}\!\vec{x}_{\mathrm{s}}^{(i)}=\vec{f}\big(\vec{x}_{\mathrm{s}}^{(i)},t\big)\overleftarrow{\mathrm{d}}\!t+\mathbf{\Sigma}^{1/2}\overleftarrow{\mathrm{d}}\!\vec{B}^{(i)}+\mathbf{\Sigma}(\mathbf{C}_{\mathrm{loc,2}}\circ\mathbf{P}_{\mathrm{f}})^{-1}\big(\vec{x}_{\mathrm{s}}^{(i)}-\vec{x}_{\mathrm{f}}^{(i)}\big)\overleftarrow{\mathrm{d}}\!t.$ (14)
Inflation. With localization, sampling error and spurious correlations are reduced, but still may accumulate over time. A practical solution to this is multiplicative inflation, which rescales ensemble anomalies around the mean by a factor $\delta^{2}>1$ (Anderson and Anderson, 1999):
$\vec{x}^{(i)}\longleftarrow\overline{\vec{x}}+\delta\big(\vec{x}^{(i)}-\overline{\vec{x}}\big),\qquad i=1,\dots,m.$ (15)
The multiplicative inflation Eq. (15) aims to compensate for sampling errors that are the indirect consequence of the nonlinearity in the ensemble forecast, which helps to cure the intrinsic error of ensemble filter schemes (Asch et al., 2016). In addition, inflation should be applied only for the forward filtering, not the backward smoothing pass. Since sampling errors are already accounted for in the filter, applying inflation in the backward update would result in a suboptimal solution (Cosme et al., 2010).